# MAPLE: Self-supervised Learning-Enhanced Nonlinear Dimensionality Reduction for Visual Analysis

### v1

@@ IV-D Parameter Sensitivity Analysis
To assess the robustness of MAPLE with respect to its two primary hyperparameters—the neighborhood size ($k$) and the trade-off parameter ($\lambda$) we conducted a grid search analysis on the Fashion-MNIST dataset (N=40, [^0] N=40{,}000). The search space included $k\in\{5,15,30\}$ and $\lambda\in\{0,0.1,0.3,0.5,0.7,0.9,2.0\}$. We evaluated the resulting layouts using our two core quantitative measures, neighborhood hit and kkNN classification accuracy, to examine whether MAPLE requires careful hyperparameter tuning or offers a broad, stable behavior. The results are summarized in Fig. 6.
@@
In contrast, the neighborhood size $k$ plays a more influential role. MAPLE achieves its highest quantitative scores at smaller neighborhood sizes ($k=5$), as neighborhood hit is a highly local metric. This behavior is also reflected qualitatively in Fig. 6(c), where structurally similar classes such as ‘Boot’, ‘Sandal’, and ‘Sneaker’ (the blue–brown–grey cluster on the left) appear more clearly separated than at $k=30$. We hypothesize that smaller $k$ values yield smaller and more homogeneous neighborhood “views”, which are better aligned with the locality assumptions in MMCR-style self-supervised learning [yerxa2023mmcr, isik2023information]. To contextualize these results, we further compare MAPLE against UMAP using matched neighborhood sizes; detailed neighborhood hit scores for each method are reported in the supplementary materials. MAPLE provides neighborhood improvements at smaller $k$, while larger $k$ values yield stronger relative gains. During development, we also explored soft probabilistic kkNN assignments, but observed only marginal improvements over the standard formulation. This suggests that binary selection of high-confidence neighbors is sufficient for MAPLE to learn meaningful manifold structure, offering a favorable balance between simplicity and accuracy.
@@ V Applying MAPLE to Single-cell Datasets
Data exploration in C. elegans embryogenesis.  We selected the C. elegans dataset because every individual worm develops through the same sequence of cell divisions, and its cell lineage has been largely mapped. These characteristics make it ideal for closely examining the DR layouts to see how they can help describe gene expression patterns and trace transcriptome dynamics during differentiation and development. The dataset we use is from Parker et al. [packer2019lineage], which contains 86,204 cells with 20,222 measured gene expression dimensions, spanning a wide diversity of cell types. It is fully annotated, and many subsequent studies [fung2020cell, sivaramakrishnan2023transcript] have validated the accuracy of these annotations. For each cell, annotations are provided at both broad levels (e.g., neurons and glia) and finer subtypes (e.g., amphid sensory neurons such as ADF, AWB), together with lineage information such as “parent” or “neuroblast” states, and rich metadata for analysis. These annotations facilitate comparison of proximity and expression differences across multiple scales of analysis.
@@
MAPLE brings better cell subtype disambiguation.  During the evaluation, the expert confirmed that the separations produced by MAPLE align mostly with the known C. elegans cell lineage. For example, as shown in Fig. 7a, MAPLE provides clearer separation of both embryo time and known cell subtypes compared to UMAP. Amphid sensory neurons such as ADF and AWC, which often appear clustered together in UMAP, are also resolved into distinct trajectories in MAPLE. At the same time, MAPLE preserves alignments between closely related categories such as ADF and their progenitor ADF_AWB, which are consistent with developmental biology findings [packer2019lineage]. This combination of proper separation and alignment with biological events gives domain experts confidence that MAPLE’s higher-resolution results more faithfully capture lineage structure than UMAP.
MAPLE facilitates hypothesis generation.  Beyond validation, MAPLE also supports the discovery of potential new insights. As illustrated in Fig. 7b, MAPLE not only disambiguates progenitor states such as ASI_parent and ASK_parent, but also highlights individual cells that act as “bridges” between trajectories [packer2019lineage]. For instance, a cell that UMAP merges into the AUA cluster is instead positioned by MAPLE between ASJ and AUA lineages. Such bridge-like cells could represent transitional populations where progenitors begin differentiating, which can be captured through the expression of marker genes of both descendant types [wagner_lineage_2020]. Identifying these rare states is valuable because it enables researchers to generate and test hypotheses about the molecular signatures of ancestor cells and the order in which lineages branch. Those insights would remain hidden in other DR layouts.
MAPLE can be used as a stand-alone method, or as a supplementary panel to mitigate the “hallucination” of connectivity that has been reported for UMAP [sun2019accuracy]. While UMAP is widely used in single-cell studies, it has also been criticized for generating overly confident continuous trajectories that may not exist [sun2019accuracy, xia2024statistical, chari2023specious]. This can mislead analysts into forming incorrect hypotheses, especially in datasets where annotations are lacking and uncertainty is high. MAPLE’s local compression and diversification strategy produces clearer separations, which may raise new questions about connectivity, as well as providing additional perspective to help validate or challenge the results of other DR methods.
@@ VI Discussion and Future Work
Conceptually, MAPLE belongs to a growing line of work that focuses on the often overlooked yet critical aspect of input-space modeling in DR methods [bohm2022unsupervised, wang2025dimension]. By reinterpreting neighborhood graphs as implicit views, MAPLE establishes a connection between MVSSL and DR, while introducing geometry-aware objectives to regularize local structure. At the same time, MAPLE represents a hybrid approach, uniquely combining parametric graph construction with non-parametric projection. We believe this opens up broader discussions on the multifaceted relationships between DR and SSL, as discussed in prior theoretical and empirical studies [hu2023your, damrich2022tsne].
Furthermore, MAPLE highlights the need to reconsider how DR methods are evaluated. For methods like ours, quantitative evaluation should move beyond statistical consistency checks between input and output space and instead develop new measures—potentially inspired by computer vision and human perception [taha2015metrics, zhang2018unreasonable]—that can better capture downstream, task-oriented goals. In addition, approaches developed in medical image segmentation [taha2015metrics], where evaluation is designed to quantitatively reflect perceptual quality and task performance, could potentially be adapted to expand the current set of evaluation measures in MAPLE.

### v2

@@ IV-D Parameter Sensitivity Analysis
To assess the robustness of MAPLE with respect to its two primary hyperparameters, the neighborhood size ($k$) and the trade-off parameter ($\lambda$), we conducted a grid search analysis on the Fashion-MNIST dataset (N=40, [^0] N=40{,}000). The search space included $k\in\{5,15,30\}$ and $\lambda\in\{0,0.1,0.3,0.5,0.7,0.9,2.0\}$. We evaluated the resulting layouts using our two core quantitative measures, neighborhood hit and kkNN classification accuracy, to examine whether MAPLE requires careful hyperparameter tuning or offers a broad, stable behavior. The results are summarized in Fig. 6, where different line styles (e.g., dashed) denote neighborhood sizes $k\in\{5,15,30\}$.
@@
In contrast, the neighborhood size $k$ plays a more influential role. MAPLE achieves its highest quantitative scores at smaller neighborhood sizes ($k=5$), as neighborhood hit is a highly local metric. This behavior is also reflected qualitatively in Fig. 6(c), where structurally similar classes such as ‘Ankle Boot’, ‘Sandal’, and ‘Sneaker’ (the blue–brown–grey cluster on the left) appear more clearly separated than at $k=30$. We hypothesize that smaller $k$ values yield smaller and more homogeneous neighborhood “views”, which are better aligned with the locality assumptions in MMCR-style self-supervised learning [yerxa2023mmcr, isik2023information]. To contextualize these results, we further compare MAPLE against UMAP using matched neighborhood sizes; detailed neighborhood hit scores for each method are reported in the supplementary materials. MAPLE provides neighborhood improvements at smaller $k$, while larger $k$ values yield stronger relative gains. During development, we also explored soft probabilistic kkNN assignments, but observed only marginal improvements over the standard formulation. This suggests that binary selection of high-confidence neighbors is sufficient for MAPLE to learn meaningful manifold structure, offering a favorable balance between simplicity and accuracy.
@@ V Applying MAPLE to Single-cell Datasets
Data exploration in C. elegans embryogenesis.  We selected the C. elegans dataset because every individual worm develops through the same sequence of cell divisions, and its cell lineage has been largely mapped. These characteristics make it ideal for closely examining the DR layouts to see how they can help describe gene expression patterns and trace transcriptome dynamics during differentiation and development. The dataset we use is from Packer et al. [packer2019lineage], which contains 86,204 cells with 20,222 measured gene expression dimensions, spanning a wide diversity of cell types. It is fully annotated, and many subsequent studies [fung2020cell, sivaramakrishnan2023transcript] have validated the accuracy of these annotations. For each cell, annotations are provided at both broad levels (e.g., neurons and glia) and finer subtypes (e.g., amphid sensory neurons such as ADF, AWB), together with lineage information such as “parent” or “neuroblast” states, and rich metadata for analysis. These annotations facilitate comparison of proximity and expression differences across multiple scales of analysis.
@@
MAPLE brings better cell subtype disambiguation.  During the evaluation, the expert confirmed that the separations produced by MAPLE align mostly with the known C. elegans cell lineage. For example, as shown in Fig. 7(a), MAPLE provides clearer separation of both embryo time and known cell subtypes compared to UMAP. Amphid sensory neurons such as ADF and AWC, which often appear clustered together in UMAP, are also resolved into distinct trajectories in MAPLE. At the same time, MAPLE preserves alignments between closely related categories such as ADF and their progenitor ADF_AWB, which are consistent with developmental biology findings [packer2019lineage]. This combination of proper separation and alignment with biological events gives domain experts confidence that MAPLE’s higher-resolution results more faithfully capture lineage structure than UMAP.
MAPLE facilitates hypothesis generation.  Beyond validation, MAPLE also supports the discovery of potential new insights. As illustrated in Fig. 7(b), MAPLE not only disambiguates progenitor states such as ASI_parent and ASK_parent, but also highlights individual cells that act as “bridges” between trajectories [packer2019lineage]. For instance, a cell that UMAP merges into the AUA cluster is instead positioned by MAPLE between ASJ and AUA lineages. Such bridge-like cells may represent transitional populations where progenitors begin differentiating, which may be captured through the expression of marker genes of both descendant types [wagner_lineage_2020]. Identifying these rare states is valuable because it enables researchers to generate and test hypotheses about the molecular signatures of ancestor cells and the order in which lineages branch. Those insights would remain hidden in other DR layouts.
MAPLE can be used as a stand-alone method, or as a supplementary panel to mitigate the “hallucination” of connectivity that has been reported for UMAP [sun2019accuracy]. While UMAP is widely used in single-cell studies, it has also been criticized for generating overly confident continuous trajectories that may not exist [sun2019accuracy, xia2024statistical, chari2023specious]. This can mislead analysts into forming incorrect hypotheses, especially in datasets where annotations are lacking and uncertainty is high. MAPLE’s local compression and diversification strategy produces clearer separations, which may raise new questions about connectivity while also providing additional perspective to help validate or challenge the results of other DR methods.
@@ VI Discussion and Future Work
Conceptually, MAPLE belongs to a growing line of work that focuses on the often-overlooked yet critical aspect of input-space modeling in DR methods [bohm2022unsupervised, wang2025dimension]. By reinterpreting neighborhood graphs as implicit views, MAPLE establishes a connection between MVSSL and DR, while introducing geometry-aware objectives to regularize local structure. At the same time, MAPLE represents a hybrid approach, uniquely combining parametric graph construction with non-parametric projection. We believe this opens up broader discussions on the multifaceted relationships between DR and SSL, as discussed in prior theoretical and empirical studies [hu2023your, damrich2022tsne].
Furthermore, MAPLE highlights the need to reconsider how DR methods are evaluated. For methods like ours, quantitative evaluation should move beyond statistical consistency checks between input and output space and instead develop new measures—potentially inspired by computer vision and human perception [taha2015metrics, zhang2018unreasonable, bae2025uncovering]—that can better capture downstream, task-oriented goals. In addition, approaches developed in medical image segmentation [taha2015metrics], where evaluation is designed to quantitatively reflect perceptual quality and task performance, could potentially be adapted to expand the current set of evaluation measures in MAPLE.